OpenStudy (anonymous):

Given log 3=0.4471, log 5=0.699, log7=0.8451, log2=0.3010, find the value of

OpenStudy (anonymous):

\[\log \frac{ \sqrt[3]{245} }{ 315 }\]

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Oooh... Prime factorisation lang katapat niyan. at wag maarte, pang grade-3 yun

terenzreignz (terenzreignz):

\[\Large \frac{\sqrt[3]{245}}{315}= \frac{\sqrt[3]{5\cdot 7^2}}{3^2\cdot 5\cdot 7}\]

OpenStudy (anonymous):

nag sosolve sa papel eh, -____-

terenzreignz (terenzreignz):

bagal eh :P

OpenStudy (anonymous):

nawawala pen ko ehh

OpenStudy (anonymous):



terenzreignz (terenzreignz):

wtf

OpenStudy (anonymous):



terenzreignz (terenzreignz):

binigay ko nga sayo eh

terenzreignz (terenzreignz):

at hindi basta basta nakaka cancel nakalilimutan mo yata... \[\Large \sqrt[3]{abc}= a^\frac13b^\frac13c^\frac13\]

OpenStudy (anonymous):

aiyt,,, ahahah

OpenStudy (anonymous):

ano sunod?

terenzreignz (terenzreignz):

Well, duh...\[\Large \log\left(\frac{\sqrt[3]{245}}{315}\right)= \log\left(\frac{\sqrt[3]{5\cdot 7^2}}{3^2\cdot 5\cdot 7}\right)\] I mean, geez, do I have to do everything? XD

OpenStudy (anonymous):

yup, xDD

OpenStudy (anonymous):

\[\log \frac{ \sqrt[3]{245} }{315 }=\log \left( 245 \right)^{\frac{ 1 }{ 3 }}-\log315\] \[=\frac{ 1 }{ 3 }\log \left( 5*7^{2} \right)-\log \left( 5*7*3^{2} \right)\] \[=\frac{ 1 }{ 3 }\log 5+\frac{ 2 }{3 }\log7-\log 5-\log7-2\log3\] \[=-\frac{ 2 }{3 }\log5-\frac{ 1 }{3 }\log7-2\log 3\] sbstitute the values and get the solution.

terenzreignz (terenzreignz):

Not forgetting, of course, that \[\Large \log(\sqrt[3]M) = \log(M^\frac13) = \frac13\log(M) \]

OpenStudy (anonymous):

ahah,xD

OpenStudy (anonymous):

yeah,, i understand it now, xDD

terenzreignz (terenzreignz):

The rest is up to you.... I'm sure you can do this :P If not, make a bloody peanut-butter sandwich and try again... it seems to be the solution to everything :D

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\log(5.7)-\log(3^2.5.7)\]

OpenStudy (anonymous):

T_T

terenzreignz (terenzreignz):

<whitling sounds...> \[\huge \frac{ 1 }{ 3 }\log(5.7^{\color{red}2 })-\log(3^2.5.7)\]

OpenStudy (anonymous):

sorry,, didn't noticed that

terenzreignz (terenzreignz):

I'm sure :P Confident with logs now?